Determine any extreme points to 

?(?) = (?-3)2 ?x in the interval [0, 4].
Answer #1

Arr: f(x) = (x-2)*** For extreme points, f(x) =o F(x) = 2(x-3)* + (x-3)ex (x-3) et X- 21 so, x = 3, 1, So, at x = 3, f(3)

If the answer helped please upvote it means a lot. For any query please comment.